# Can LLMs Understand Time Series Anomalies?

### v2

@@ Abstract
Large Language Models (LLMs) have gained popularity in time series forecasting, but their potential for anomaly detection remains largely unexplored. Our study investigates whether LLMs can understand and detect anomalies in time series data, focusing on zero-shot and few-shot scenarios. Inspired by conjectures about LLMs’ behavior from time series forecasting research, we formulate key hypotheses about LLMs’ capabilities in time series anomaly detection. We design and conduct principled experiments to test each of these hypotheses. Our investigation reveals several surprising findings about LLMs for time series: (1) LLMs understand time series better as images rather than as text, (2) LLMs did not demonstrate enhanced performance when prompted to engage in explicit reasoning about time series analysis. (3) Contrary to common beliefs, LLM’s understanding of time series do not stem from their repetition biases or arithmetic abilities. (4) LLMs’ behaviors and performance in time series analysis vary significantly across different model architectures. This study provides the first comprehensive analysis of contemporary LLM capabilities in time series anomaly detection. Our results suggest that while LLMs can understand trivial time series anomalies (we have no evidence that they can understand more subtle real-world anomalies), many common conjectures based on their reasoning capabilities do not hold. Our code and data are available at https://github.com/Rose-STL-Lab/AnomLLM/.
@@ 1 Introduction
The remarkable progress in large language models (LLMs) has led to their application in various domains, including time series analysis. Recent studies have demonstrated LLMs’ potential as zero-shot and few-shot learners in tasks such as forecasting and classification (Gruver et al., 2023; Liu et al., 2024c). However, the effectiveness of LLMs in time series analysis remains a subject of debate. While some researchers argue that LLMs can leverage their pretrained knowledge to understand time series patterns (Gruver et al., 2023), others suggest that simpler models may outperform LLM-based approaches (Tan et al., 2024).
This debate raises a fundamental question: Do LLMs truly understand time series? To address this question, we must consider not only the models’ predictive performance but also their ability to recognize normal patterns, identify abnormal behavior, and provide explanations for their decisions. This understanding requires a deeper understanding of the underlying temporal dynamics.
In this paper, we present the first comprehensive investigation into LLMs’ understanding of time series data through the lens of anomaly detection. We focus on the behavior of state-of-the-art LLMs and multimodal LLMs (M-LLMs) across different anomaly types under controlled conditions. Our evaluation strategy incorporates multimodal inputs (textual and visual representations of time series), various prompting techniques, and structured output formats. The results are quantitatively assessed using the affinity F1 score to challenge unvalidated conjectures and claims about LLMs’ time series understanding in prior works. This work leads to a more nuanced understanding of LLMs’ capabilities and limitations in processing time series data.
We provide empirical evidence that contradicts several prevailing beliefs about LLMs’ abilities in time series analysis, revealing:
  • Visual Advantage: LLMs perform significantly better when processing visualized time series data compared to textual representations.
  • Limited Reasoning: Contrary to expectations, LLMs do not extensively leverage complex reasoning processes when analyzing time series data. Their performance often decreases when prompted to explain their reasoning.
  • Non-Human-Like Processing: LLMs’ approach to anomaly detection differs from human perception. They can identify subtle trends that humans might miss and their performance is not tied to arithmetic abilities.
  • Model-Specific Capabilities: Time series understanding and anomaly detection capabilities differ across various LLM architectures, highlighting the importance of model selection.
@@ 2 Related Work
LLMs have been applied to various time series analysis tasks, including forecasting and classification. These models have shown promising results in capturing temporal dependencies and patterns in time series data. Gruver et al. (2023) demonstrated that LLMs like GPT-3 and LLaMA-2 can perform zero-shot time series forecasting by encoding time series as strings of numerical digits, achieving comparable performance to purpose-built models. Liu et al. (2024c) proposed a Cross-Modal LLM Fine-Tuning framework to address the distribution discrepancy between textual and temporal input tokens in LLM-based multivariate time series forecasting. Liu et al. (2024b) introduced Time-MMD, a multi-domain, multimodal time series dataset for LLM finetuning. For anomaly detection, Liu et al. (2024a) proposed AnomalyLLM, a knowledge distillation-based approach using GPT-2 as the teacher network. Zhang et al. (2024) provided a comprehensive survey of LLM applications in time series analysis. In the financial domain, Wimmer & Rekabsaz (2023) used vision language models, i.e., CLIP, but not M-LLMs to process image representations of stock data for market change prediction.
However, the effectiveness of LLMs for time series analysis remains controversial. Zeng et al. (2022) argued that the permutation-invariant nature of self-attention mechanisms may lead to loss of critical temporal information. Tan et al. (2024) found that removing the LLM component or replacing it with a basic attention layer often improved performance in popular LLM-based forecasting methods. The interpretability of LLMs in time series analysis remains a challenge, as their reasoning capabilities are often opaque and difficult to interpret.
@@
Time series anomaly detection is a critical task in various domains, including finance, healthcare, and cybersecurity. Traditional methods rely on statistical techniques, while recent work has focused on developing deep learning-based approaches (Audibert et al., 2022; Chen et al., 2022; Tuli et al., 2022). Audibert et al. (2022) compared conventional, machine learning-based, and deep neural network methods, finding that no family of methods consistently outperforms the others. Chen et al. (2022) proposed a deep variational graph convolutional recurrent network for multivariate time series anomaly detection. Tuli et al. (2022) introduced a transformer-based anomaly detection model with adversarial training and meta-learning.
@@
One reason for using deep learning models in time series anomaly detection is their ability to bring prior knowledge on what constitutes normal behavior from pretraining on large-scale datasets. Large language models (LLMs) may offer a promising solution due to their strong zero-shot capabilities. However, there is currently a lack of research exploring the application of modern (M-)LLMs to time series anomaly detection, presenting an opportunity for investigation in this area.
@@
The multimodal capabilities of LLMs have been explored in various domains, including image captioning, video understanding, and multimodal translation (Lu et al., 2019; Li et al., 2019; Sun et al., 2019; Huang et al., 2019). Recent advancements have led to more sophisticated M-LLMs, such as Qwen-VL and Phi-3-Vision, demonstrating superior performance in visual-centric tasks and compact deployment capabilities (Bai et al., 2023; Abdin et al., 2024). In the context of time series analysis, M-LLMs have been used to model multimodal data, such as time series and textual information, showing promising results in forecasting and anomaly detection (Liu et al., 2021). However, there is a notable lack of application of M-LLMs to time series as visual inputs, even though humans often detect time series anomalies through visual inspection. This gap is particularly significant given that time series data can be represented in multiple modalities (e.g., numerical, textual, or visual) without losing substantial new information. Consequently, time series analysis presents a unique opportunity to evaluate an M-LLM’s ability to understand and process the same underlying data across different representational formats. The potential for M-LLMs to bridge the gap between human visual intuition and machine learning in time series analysis represents a promising avenue for future research.
@@ 3 Time Series Anomaly Detection: Definition and Categorization
We begin by defining time series anomaly detection and categorizing different anomalies.
3.1 Time Series Anomaly Detection Definition
We consider time series $X:=\{x_{1},x_{2},\ldots,x_{T}\}$ collected at regular intervals, where $x_{t}$ is the feature scalar or vector at time $t$, and $T$ is the total number of time points. Anomalies are data points that deviate significantly from the expected pattern of the time series and can be defined by the generating function or conditional probability, depending on whether the system is deterministic or stochastic.
@@ 3.1 Time Series Anomaly Detection Definition
Interval-Based Anomalies.
It is common to define anomalies as continuous intervals of time points rather than individual points. Let $R$ be a set of time intervals where anomalies occur, defined as $R=\{[t_{start}^{1},t_{end}^{1}],[t_{start}^{2},t_{end}^{2}],\ldots,[t_{start}^{k},t_{end}^{k}]\}$, where $[t_{start}^{i},t_{end}^{i}]$ represents the $i$-th anomalous time interval, with $t_{start}^{i}$ and $t_{end}^{i}$ being the start and end times of the interval, respectively. The binary labels for this definition would be:
$y_{t}=\begin{cases}1&\text{if }t\in[t_{start}^{i},t_{end}^{i}]\text{ for all }i\in\{1,\ldots,k\}\\ 0&\text{otherwise}\end{cases}$
We will consider this definition of anomalies in our evaluation of LLMs’ anomaly detection capabilities. Note that we are interested in detecting anomalies at the data-point level instead of the sequence level. While we use interval-based anomalies, these are still composed of individual anomalous data points within a sequence. This approach differs from sequence-level anomaly detection, where entire sequences would be classified as anomalous or normal.
@@
3.2 Time Series Anomalies Categorization
Figure 1: Example time series with different anomaly types
Time series anomalies can be broadly categorized into two main types: out-of-range anomalies and contextual anomalies (Lai et al., 2024). The contextual anomalies can be further divided into frequency anomalies, trend anomalies, and contextual point anomalies. Each type presents unique characteristics and challenges for detection. By examining how LLMs recognize different types of anomalies, we can demonstrate that our hypotheses about LLMs’ understanding of time series data generalize across anomaly types.
@@ 3.2.1 Out-of-range Anomalies
Out-of-range anomalies are individual data points far from all other data points in the series such that they can be detected even if the time series is shuffled, see Figure 1(d). These anomalies can be identified using non-temporal methods, such as clustering or thresholding.
@@ 3.2.2 Contextual Anomalies
Trend Anomalies.  Trend anomalies occur when the overall direction or rate of change in a time series deviates significantly from the expected trend. These anomalies are characterized by unexpected changes in the changing rate of the time series. A trend anomaly may manifest as a sudden acceleration, deceleration, or reversal of the established trend. Figure 1(a) shows an example of acceleration. The detection of trend anomalies often involves inspection of the gradient.
@@
As most of the literature on LLM for time series analysis focuses on forecasting, we use the conjectures from these works as a starting point to understand LLMs’ behavior in anomaly detection. The tasks of time series forecasting and anomaly detection share many similarities. By definition, deterministic forecasting of future time steps is about learning the generating function (see Equation 1). Probabilistic forecasting is about learning the conditional probability function (see Equation 2). Therefore, both time series forecasting and anomaly detection rely heavily on extrapolation. In forecasting tasks, models extrapolate past patterns to predict future values, extending the known series into unknown territory. Similarly, anomaly detection involves extrapolating the “normal” behavior of a time series to identify points that deviate significantly from this expected pattern.
LLMs are trained on a corpus of token sequences, $\{U_{1},U_{2},\ldots,U_{N}\}$, where $U_{i}=\{u_{1},u_{2},\ldots,u_{L_{i}}\}$ and $u_{j}$ is a token in the vocabulary $\mathcal{V}$. The model learns to autoregressively predict the next token in the sequence given the previous tokens, i.e., $P(u_{j+1}|u_{1},u_{2},\ldots,u_{j})$. The motivation for applying LLMs to time series forecasting is often their zero-shot extrapolation capabilities (Brown et al., 2020). The autoregressive generation of tokens and time series steps (in Equation 2) are similar, and the LLMs act as an Occam’s razor to find the simplest form of $G$ (in Equation 1) (Gruver et al., 2023). This connection suggests that hypotheses made in LLMs for forecasting may also apply to anomaly detection. Many such hypotheses are proposed as possible explanations for the model’s behavior without validation from controlled studies, which motivates our investigation.
@@ 4.1 Hypotheses
Hypothesis 1 (Tan et al., 2024) on CoT Reasoning LLMs do not benefit from engaging in step-by-step reasoning about time series data.
The authors claim that existing LLM methods, including zero-shot approaches, do very little to use innate reasoning. While they demonstrate that LLM-based methods perform similarly or worse than methods without LLMs in time series tasks, none of their experiments assess the LLMs’ reasoning capabilities.
To validate this hypothesis, we focus on the performance of one-shot and zero-shot Chain of Thought (CoT) prompting, which explicitly elicits an LLM’s reasoning abilities (Wei et al., 2022). We use terms from cognitive science to describe the LLMs’ different behaviors with and without CoT: the reflexive mode (slow, deliberate, and logical) and the reflective mode (fast, intuitive, and emotional) (Lieberman, 2003). Therefore, the question becomes whether the LLMs benefit from the reflexive mode, when it thinks slowly about the time series. If the hypothesis is false, the LLMs should perform better when they are prompted to explain.
Hypothesis 2 (Gruver et al., 2023) on Repetition Bias LLMs’ repetition bias (Holtzman et al., 2020) corresponds precisely to its ability to identify and extrapolate periodic structure in the time series.
This hypothesis draws a parallel between LLMs’ tendency to generate repetitive tokens and their potential ability to recognize periodic patterns in time series data. To validate the hypothesis, we design an experiment where the datasets contain both perfectly periodic and noisy periodic time series. The introduction of minor noise would disrupt the exact repetition of tokens in the input sequence, even if the underlying pattern remains numerically approximately periodic. If the hypothesis holds, we should observe a significant drop in performance, despite the numbers maintaining its fundamental periodic structure.
@@
This hypothesis suggests a connection between LLMs’ arithmetic capabilities and their ability to extrapolate simple mathematical sequences. It is argued that LLM’s proficiency in basic arithmetic operations, such as addition and multiplication, enables it to extend patterns like linear sequences (e.g., $x(t)=2t$) by iteratively applying the addition operation (e.g., $+2$). To validate the hypothesis, we design an experiment where an LLM is specifically guided to impair its arithmetic abilities while preserving its other linguistic and reasoning capabilities. If the hypothesis holds, we should observe a corresponding decline in the model’s ability to predict anomalies in the trend datasets. Otherwise, LLMs rely on alternative mechanisms for time series pattern recognition and extrapolation.
@@
Hypothesis 4 on Visual Reasoning Time series anomalies can be more easily detected as visual input rather than text input.
@@
Motivated by the fact that human analysts often rely on visual representations for anomaly detection in time series, we hypothesize that M-LLMs, whose training data includes human expert detection tasks, may prefer visualized time series to raw numerical data. This hypothesis can be readily tested by comparing the performance of multimodal LLMs on identical time series presented as both text and visualizations. A significant performance gap would suggest LLMs, like humans, prefer to understand time series data visually.
Hypothesis 5 on Visual Perception Bias LLMs’ understanding of anomalies is consistent with human visual perception.
@@
We hypotheses based on the growing interest in using LLMs for time series due to their internal human-like knowledge (Jin et al., 2024b) and their demonstrated ability to align with human cognition in complex tasks (Thomas et al., 2024). However, humans have known cognitive limitations in detecting certain types of anomalies, and recent research suggests that LLMs may exhibit human-like cognitive biases (Opedal et al., 2024).
@@
To validate this hypothesis, we leverage findings from human perception research, such as the observation that humans are far less sensitive to acceleration than to constant speed changes (Mueller & Timney, 2016). We compare LLM performance on two datasets: one featuring anomalies that revert an increasing trend to a decreasing trend (analogous to negating constant speed), and another with anomalies that accelerate an increasing trend. Both datasets would have similar prevalence rates, making them equally detectable by traditional methods that identify gradient changes. If LLMs exhibit significantly poorer performance in detecting acceleration anomalies compared to trend reversals, it would suggest that they indeed suffer from similar perceptual limitations as humans.
@@
Hypothesis 6 on Long Context Bias LLMs perform better with time series with fewer tokens, even if there is information loss.
@@
Hypothesis 7 on Architecture Bias LLMs’ time series understanding are consistent across different model architectures.
This hypothesis stems from the tendency in recent literature (Tang et al., 2024; Tan et al., 2024; Zeng et al., 2022) to generalize findings about time series understanding across all LLMs based on experiments with a limited set of models, typically GPT and LLaMA variants. This approach implies that different model families’ comprehension of time series data is universally consistent, varying primarily with the number of parameters. However, unlike in NLP tasks where specific models excel in areas like translation or mathematics (Cobbe et al., 2021), there’s a lack of understanding regarding specialized skills in time series analysis across different LLM architectures. To test this hypothesis, we perform all previous experiments across different LLM architectures. If the hypothesis holds, we should see previous conclusions either consistently validated or invalidated across all models.

### v3

@@ Abstract
Large Language Models (LLMs) have gained popularity in time series forecasting, but their potential for anomaly detection remains largely unexplored. Our study investigates whether LLMs can understand and detect anomalies in time series data, focusing on zero-shot and few-shot scenarios. Inspired by conjectures about LLMs’ behavior from time series forecasting research, we formulate key hypotheses about LLMs’ capabilities in time series anomaly detection. We design and conduct principled experiments to test each of these hypotheses. Our investigation reveals several surprising findings about LLMs for time series: (1) LLMs understand time series better as images rather than as text, (2) LLMs do not demonstrate enhanced performance when prompted to engage in explicit reasoning about time series analysis. (3) Contrary to common beliefs, LLMs’ understanding of time series does not stem from their repetition biases or arithmetic abilities. (4) LLMs’ behaviors and performance in time series analysis vary significantly across different models. This study provides the first comprehensive analysis of contemporary LLM capabilities in time series anomaly detection. Our results suggest that while LLMs can understand trivial time series anomalies, we have no evidence that they can understand more subtle real-world anomalies. Many common conjectures based on their reasoning capabilities do not hold. All synthetic dataset generators, final prompts, and evaluation scripts have been made available in https://github.com/rose-stl-lab/anomllm.
@@ 1 Introduction
The remarkable progress in large language models (LLMs) has led to their application in various domains, including time series analysis. Recent studies have demonstrated LLMs’ potential as zero-shot and few-shot learners in tasks such as forecasting and classification (Gruver et al., 2023; Liu et al., 2024c). However, the effectiveness of LLMs in time series analysis remains a subject of debate. While some researchers argue that LLMs can leverage their pretrained knowledge to understand time series patterns (Gruver et al., 2023), others suggest that simpler models can match or outperform LLMs (Tan et al., 2024).
This debate raises a fundamental question: Do LLMs truly understand time series? To address this question, we must look beyond models’ predictive performance. Forecasting typically hinges on metrics like MSE, which can overlook deeper model understanding. A model that simply outputs a nearly constant line might still achieve a passable MSE but reveals little about its capacity to interpret dynamics. Shifting our focus to anomaly detection changes the game: it forces LLMs to pinpoint irregular behavior and thus tests whether they grasp the underlying patterns, not just how well they extrapolate an average.
In this paper, we present the first comprehensive investigation into LLMs’ understanding of time series data through the lens of anomaly detection. We focus on state-of-the-art LLMs and multimodal LLMs (M-LLMs) across different anomaly types under controlled conditions. Our evaluation strategy incorporates multimodal inputs (textual and visual representations of time series), various prompting techniques, and structured output formats. The results are quantitatively assessed using the affinity F1 score. We provide empirical evidence to challenge existing conjectures and claims about LLMs’ time series understanding. Our findings reveal a more nuanced understanding of LLMs’ capabilities and limitations in time series data, including:
  • Visual Advantage: LLMs perform significantly better in anomaly detection when processing time series data as images rather than text tokens.
  • Limited Reasoning: LLMs do not benefit from chain-of-thought reasoning when analyzing time series data. Their performance often decreases when prompted to explain their reasoning.
  • Alternative Processing Mechanisms: Contrary to common beliefs, LLMs’ understanding of time series does not stem from their repetition biases or arithmetic abilities, challenging prevailing assumptions about how these models process numerical data.
  • Model Heterogeneity: Time series understanding and anomaly detection capabilities differ across various LLM architectures, highlighting the importance of model selection.
In summary, while LLMs can detect simple anomalies, their abilities to understand and reason about numerical time series are considerably limited. Our research calls for more caution in applying LLMs to time series and for controlled studies to examine LLMs’ behavior for other data modalities.
@@ 2 Related Work
LLMs have been applied to various time series analysis tasks, with recent literature establishing strong claims about their capabilities. Gruver et al. (2023) showed that LLMs such as GPT-3 and LLaMA-2 possess broad pattern extrapolation capabilities, enabling zero-shot time series forecasting by encoding time series as strings of numerical digits and achieving comparable performance to purpose-built models. Building on these claims, Liu et al. (2024c) proposed a Cross-Modal LLM Fine-Tuning framework, suggesting that LLMs can provide interpretable predictions while addressing the distribution discrepancy between textual and temporal input tokens in multivariate time series forecasting. Liu et al. (2024b) introduced Time-MMD, a multi-domain, multimodal time series dataset for LLM finetuning. For anomaly detection, Liu et al. (2024a) proposed AnomalyLLM, a knowledge distillation-based approach using GPT-2 as the teacher network. Zhang et al. (2024) provided a comprehensive survey of LLM applications in time series analysis. In the financial domain, Wimmer & Rekabsaz (2023) used vision language models, i.e., CLIP, but not M-LLMs to process visualizations of stock data for market change prediction.
However, these prevailing beliefs about LLMs’ pattern extrapolation capabilities and interpretable predictions remain controversial. Zeng et al. (2023) argued that the permutation-invariant nature of self-attention mechanisms may lead to loss of critical temporal information. Tan et al. (2024) found that removing the LLM component or replacing it with a basic attention layer often improved performance in popular LLM-based forecasting methods, challenging the assumed benefits of LLMs’ pattern recognition abilities. The interpretability of LLMs in time series analysis remains a challenge, as their reasoning capabilities are often opaque and difficult to interpret.
@@
Time series anomaly detection is a critical task in various domains, including finance, healthcare, and cybersecurity. Traditional methods rely on statistical techniques, while recent work has focused on developing deep learning-based approaches (Audibert et al., 2022; Chen et al., 2022; Tuli et al., 2022). Audibert et al. (2022) compared conventional, machine-learning-based, and deep neural network methods, finding that no family of methods consistently outperforms the others. Chen et al. (2022) proposed a deep variational graph convolutional recurrent network for multivariate time series anomaly detection. Tuli et al. (2022) introduced a transformer-based anomaly detection model with adversarial training and meta-learning.
@@
One reason for using deep learning models in time series anomaly detection is their ability to bring prior knowledge on what constitutes normal behavior from pretraining on large-scale datasets. Large language models (LLMs) may offer a promising solution due to their strong zero-shot capabilities. However, there is currently a lack of systematic study of modern (M-)LLMs for time series anomaly detection, which we aim to address in this work.
@@
Multimodal LLMs combine text with other data modalities and have been explored in various domains, including image captioning, video understanding, and multimodal translation (Lu et al., 2019; Li et al., 2019; Sun et al., 2019; Huang et al., 2019). Recent advancements have led to more sophisticated M-LLMs, such as Qwen-VL and Phi-3-Vision, demonstrating superior performance in visual-centric tasks and compact deployment capabilities (Bai et al., 2023; Abdin et al., 2024). In the context of time series analysis, M-LLMs have been used to model multimodal data, such as time series and textual information, showing promising results in forecasting and anomaly detection (Liu et al., 2021). However, there is a notable absence of research applying M-LLMs to time series data presented as visual inputs, even though humans often detect time series anomalies through visual inspection. This gap is particularly significant given that time series data can be represented in different modalities (e.g., numerical, textual, or visual) without losing substantial new information. Consequently, time series analysis presents a unique opportunity to evaluate M-LLMs’ ability to understand and process the same underlying data across different representational formats.
@@ 3 Time Series Anomaly Detection: Definition and Categorization
We begin by defining time series anomaly detection and categorizing different types of anomalies.
3.1 Anomaly Definition
We consider time series $X:=\{x_{1},x_{2},\ldots,x_{T}\}$ collected at regular intervals, where $x_{t}$ is the feature scalar or vector at time $t$, and $T$ is the total number of time points. Anomalies are data points that deviate significantly from the expected pattern of the time series. The expected pattern of a time series refers to the governing function or conditional probability that the data is expected to follow, depending on whether the system is deterministic or stochastic.
@@ 3.1 Anomaly Definition
Time series anomalies can be analyzed at two levels: (1) within individual sequences, where specific points or intervals deviate from the normal pattern, and (2) across different sequences, where entire sequences are considered anomalous. This paper focuses on the first level, specifically detecting anomalous intervals within individual sequences.
Anomalous Intervals.
We define anomalies as continuous intervals of time points that deviate from the expected pattern. Let $R$ be a set of anomalous time intervals: $R=\{[t_{start}^{1},t_{end}^{1}],[t_{start}^{2},t_{end}^{2}],\ldots,[t_{start}^{k},t_{end}^{k}]\}$ where $[t_{start}^{i},t_{end}^{i}]$ represents the $i$-th anomalous interval, with $t_{start}^{i}$ and $t_{end}^{i}$ being its start and end times. When $t_{start}^{i}=t_{end}^{i}$, the anomaly is a single point anomaly. For a time series $X=\{x_{1},x_{2},\ldots,x_{T}\}$, we assign binary labels:
$y_{t}=\begin{cases}1&\text{if }t\in[t_{start}^{i},t_{end}^{i}]\text{ for any }i\in\{1,\ldots,k\}\\ 0&\text{otherwise}\end{cases}$
@@
3.2 Anomaly Pattern Classification
Figure 1: Example time series with different anomaly types, with anomalous regions highlighted in red.
Patterns of time series anomalies can be categorized into two main types based on their nature: out-of-range anomalies which exceed normal value thresholds and contextual anomalies which exhibit abnormal behavior only within specific contexts (Lai et al., 2023). The contextual anomalies can be further divided into frequency anomalies, trend anomalies, and contextual point anomalies. Each type presents unique characteristics and challenges for detection. By examining how LLMs recognize these diverse anomaly types, we can verify whether our hypotheses about LLMs’ understanding of time series data hold consistently across different pattern variations.
@@ 3.2.1 Out-of-range Anomalies
Out-of-range anomalies are data points that lie far outside the normal range of values in a time series. These anomalies can be detected even when the time series order is shuffled, as shown in Figure 1(d). If a model can detect out-of-range anomalies but fails to detect contextual anomalies, this suggests it is not using the positional information in the time series (Tan et al., 2024).
@@ 3.2.2 Contextual Anomalies
Trend Anomalies.  Trend anomalies manifest as a sudden acceleration (Figure 1(a)), deceleration, or reversal of the established trend. These anomalies are characterized by unexpected changes in the changing rate of the time series, detected through gradient analysis $g_{t}=(x_{t}-x_{t-1})/\Delta t$. To reduce noise sensitivity, smoothed gradients $g_{t}^{smooth}=(MA_{t}-MA_{t-1})/\Delta t$ are often used, where $MA_{t}$ is a moving average over a window of points.
@@
As most of the literature on LLM for time series focuses on forecasting, we use the conjectures from these works as a starting point to understand LLMs’ behavior in anomaly detection. The tasks of time series forecasting and anomaly detection share many similarities. By definition, deterministic forecasting of future time steps is about learning the generating function (see Equation 1). Probabilistic forecasting is about learning the conditional probability function (see Equation 2). Therefore, both time series forecasting and anomaly detection rely heavily on extrapolation. In forecasting tasks, models extrapolate past patterns to predict future values, extending the known series into unknown territory. Similarly, anomaly detection involves extrapolating the “normal” behavior of a time series to identify points that deviate significantly from this expected pattern.
LLMs are trained on a corpus of token sequences, $\{U_{1},U_{2},\ldots,U_{N}\}$, where $U_{i}=\{u_{1},u_{2},\ldots,u_{L_{i}}\}$ and $u_{j}$ is a token in the vocabulary $\mathcal{V}$. The model learns to autoregressively predict the next token in the sequence given the previous tokens, i.e., $P(u_{j+1}|u_{1},u_{2},\ldots,u_{j})$. The motivation for applying LLMs to time series forecasting is often their zero-shot extrapolation capabilities (Brown et al., 2020). The autoregressive generation of tokens and that of time series steps (in Equation 2) are similar, and the LLMs act as an Occam’s razor to find the simplest form of $G$ (in Equation 1) (Gruver et al., 2023). This connection suggests that hypotheses made in LLMs for forecasting may also apply to anomaly detection. Many such hypotheses are proposed as possible explanations for the model’s behavior without validation from controlled studies, which motivates our investigation.
@@ 4.1 Hypotheses
Hypothesis 1 (Tan et al., 2024) on Chain-of-Thought (CoT) Reasoning LLMs do not benefit from engaging in step-by-step reasoning about time series data.
The authors claim that existing LLM methods, including zero-shot ones, do very little to use innate reasoning. While they demonstrate that LLM methods perform similarly or worse than those without LLMs in time series tasks, none of their experiments assess the LLMs’ reasoning capabilities.
To validate this hypothesis, we focus on the performance of one-shot and zero-shot CoT prompting, which explicitly elicits an LLM’s reasoning abilities (Wei et al., 2022). We use terms from cognitive science to describe the LLMs’ different behaviors with and without CoT: the reflexive mode (slow, deliberate, and logical) and the reflective mode (fast, intuitive, and emotional) (Lieberman, 2003). Therefore, the question becomes whether the LLMs benefit from the reflexive mode, when it thinks slowly about the time series. If the hypothesis is false, the LLMs should perform better when they are prompted to explain.
Hypothesis 2 (Gruver et al., 2023) on Repetition Bias LLMs’ repetition bias (Holtzman et al., 2020) corresponds precisely to their ability to identify and extrapolate periodic structure in the time series.
This hypothesis draws a parallel between LLMs’ tendency to generate repetitive tokens and their potential ability to recognize periodic patterns in time series data. To validate the hypothesis, we design an experiment where the datasets contain both perfectly periodic and noisy periodic time series. The introduction of minor noise would disrupt the exact repetition of tokens in the input sequence, even if the underlying pattern remains numerically approximately periodic. If the hypothesis holds, we should observe a significant drop in performance, despite the numbers maintaining its fundamental periodic structure. See Appendix E for formal definitions of Hypothesis 2 and 3.
@@
This hypothesis suggests a connection between LLMs’ arithmetic capabilities and their ability to extrapolate simple mathematical sequences. It is argued that LLMs’ proficiency in basic arithmetic operations, such as addition and multiplication, enables it to extend patterns like linear sequences (e.g., $x(t)=2t$) by iteratively applying the addition operation (e.g., $+2$). To validate the hypothesis, we design an experiment where an LLM is specifically guided to impair its arithmetic abilities while preserving its other linguistic and reasoning capabilities. If the hypothesis holds, we should observe a corresponding decline in the model’s ability to predict anomalies in the trend datasets. Otherwise, LLMs rely on alternative mechanisms for time series pattern recognition and extrapolation.
@@
Hypothesis 4 (Dong et al., 2024) on Visual Reasoning Time series anomalies can be more easily detected as visual input rather than text input.
@@
Motivated by the fact that human analysts often rely on visual representations for anomaly detection in time series, we hypothesize that M-LLMs, whose training data includes human expert detection tasks, may prefer time series as images rather than raw numerical data. Similar assumption is also proposed in recent work by Dong et al. (2024), who demonstrated that explicitly prompting LLMs to ”think visually” about time series improved their anomaly detection capabilities, even without actual visual input. This hypothesis can be readily tested by comparing the performance of multimodal LLMs on identical time series presented as both text and visualizations.
Hypothesis 5 on Visual Perception Bias LLMs exhibit similar detection limitations to human perceptual biases, e.g., in acceleration perception when analyzing visual time series representations.
@@
We hypothesize based on the growing interest in using LLMs due to their internal human-like knowledge (Jin et al., 2024) and their ability to align with human cognition in complex tasks (Thomas et al., 2023). However, humans have known cognitive limitations in detecting subtle anomalies, and recent research suggests that LLMs may exhibit human-like cognitive biases (Opedal et al., 2024).
@@
To validate this hypothesis, we leverage findings from human perception research. For example, Mueller & Timney (2016) reported that humans are more sensitive to sudden changes in motion direction or velocity (like trend reversals) than to gradual acceleration changes. We compare LLM performance on two datasets: one featuring anomalies that revert an increasing trend to a decreasing trend (analogous to negating constant speed), and another with anomalies that accelerate an increasing trend. Both datasets would have similar prevalence rates, making them equally detectable by traditional methods that identify gradient changes. If LLMs exhibit significantly poorer performance in detecting acceleration anomalies compared to trend reversals, it would suggest that they indeed suffer from similar perceptual limitations as humans.
@@
Hypothesis 6 on Long Context Bias LLMs perform better for time series with fewer tokens, even if there is information loss.